A puppy grows rapidly. If a fully grown labrador retriever weighs 34 kg when it reaches its full height, and your 24 week old lab puppy is 3% shorter than the average lab, how much does the puppy weigh?
1
Expert's answer
2019-01-25T12:04:07-0500

If to consider the height and the weight of fully grown labrador as 100%, then the puppy’s height and weight are 97%.

Find the mass of the puppy:

m = 34 × 0.97 = 32.98 (kg)